Circular No. 8681

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2006ao T. Boles, Coddenham, England, reports the discovery of an apparent

supernova on unfiltered CCD images taken on Mar. 1.069 (at mag 17.6) and 1.810 UT (at mag 17.4) with a 0.35-m reflector. The new object is located at α = 15^{h}59^{m}19^{s}.25, δ = +65^{o}13′43′′.6 (equinox 2000.0; average position from ten images on two nights), which is ≈ 3′′.1 east and 13′′.3 south of the center of the nearest galaxy, which itself is within a small cluster. Nothing is present at this location on Boles’ images from 2005 July 12 and May 11 (limiting mag 19.5) or on Digitized Sky Survey plates from 1994 May 31 (limiting red mag 20.5) and 1992 June 4 (limiting blue mag 21.0).

COMET 73P/SCHWASSMANN-WACHMANN D. Schleicher, Lowell Observatory, obtained narrowband photometry

of comet 73P on Feb. 25 at r = 1.65 AU using the Hall 1.1-m telescope at Lowell Observatory, with the following averaged results for component C: log Q(OH) = 27.30; equivalent log Q(water; vectorial) = 27.33; log Q(CN) = 24.89; log Q(C_{2}) = 24.01; log Afρ = 1.7 (cf. IAUC 7342). For component B: log Q(OH) = 26.8; loq Q(water) = 26.8; log Afρ = 1.1. The relative C_{2}-to-CN ratio for component C is strongly depleted (by 10 times) with respect to “normal”, as defined by A’Hearn et al. (1995, Icarus 118,

223) — consistent with the upper-limit result for 73P determined by Fink and Hicks (1996, Ap.J. 459, 729). To within relatively large uncertainties, component B exhibits the same depletion of C_{2} as does component C.

V1647 ORIONIS AND ASSOCIATED NEBULA Further to IAUC 8600 and IBVS 5661, several observers recently have

reported the extreme fading of the nebula associated with V1647 Ori, including B. Reipurth (University of Hawaii), G. M. Ratto (Pisa, Italy), and P. Corelli (Pagnacco, Italy). Corelli writes that his unfiltered CCD frames from Feb. 6 show no nebula to a limiting stellar mag of 18. Reipurth remarks that this appears to confirm its having been a short ‘EXor’ eruption (cf. IAUC 8289).

RS OPHIUCHI Additional photometry by G. Sostero and E. Guido (cf. IAUC 8673):

Feb. 21.48 UT, V = 7.36, B–V = +0.70; 22.48, 7.43, +0.66; 24.48, 7.70, +0.79; 25.48, 7.75, +0.65; 26.47, 7.76, +0.65; 27.48, 8.04, +0.62.

2006 March 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green